# Gaussian-Forest: Hierarchical-Hybrid 3D Gaussian Splatting for Compressed Scene Modeling

### v1

@@ 4.1 Implementation Details
Features and Decoders. The two MLPs, $\mathcal{F}_{\text{rgb}}$ and $\mathcal{F}_{\text{cov}}$, are implemented using the fast fully-fused-MLPs from Tiny-CUDA-NN. Each MLP consists of 2 hidden layers and is 64 neurons wide. All features are represented in 16-bit half-float, aligning with the output of the fully-fused-MLPs.
4.2 Comparative Methods
@@
3DGS Kerbl et al. (2023) stands out for its SOTA performance in rendering speed and quality, albeit with a substantial model parameter count. We primarily compare our approach with 3DGS, as we aim to preserve or even enhance the rendering speed and quality while reducing the parameter count. Additionally, we compared with three advanced ray tracing-based radiance field methods: Plenoxels Fridovich-Keil et al. (2022), based on explicit scene representation; InstantNGP Müller et al. (2022), utilizing a hybrid representation; and Mip-NeRF360 Barron et al. (2022), employing implicit representation. These three approaches represent typical examples of different scene representation methods. Contrasting with them allows for a comprehensive demonstration of the characteristics of our method.
4.3 Datasets and Metrics
@@
Following 3DGS, our model has been evaluated across 21 diverse scenarios. Of these, 13 scenes are based on real-world captures, including all nine scenes introduced by Mip-NeRF360 Barron et al. (2022), two scenes from the Tanks&Temples dataset Knapitsch et al. (2017), and two from Deep Blending Hedman et al. (2018). Additionally, all eight synthetic scenes from the Synthetic Blender dataset Mildenhall et al. (2020) are incorporated. These datasets encompass large-scale unbounded outdoor environments, indoor settings, and object-centric scenes. We employ commonly used PSNR, SSIM Wang et al. (2004), and LPIPS Zhang et al. (2018) for evaluating rendering quality. Moreover, we provide information on rendering speed, model size, and training time, along with the speed-to-size ratio for a comprehensive and straightforward comparison.
4.4 Results and Analyses
4.4.1 Quantitative Comparisons
Quantitative results across four benchmarks are presented in Table 1, accompanied by additional visualizations for quantitative comparisons on Mip-NeRF360 and Tanks&Temples showcased in Fig. 3. Firstly, our approach excels in adeptly balancing rendering speed and model size. Across all scenarios, Gaussian-Forest achieves the highest speed-to-size ratio, surpassing all other comparative methods by a large margin while ensuring high-fidelity rendering quality.
In addition, compared to the unprecedentedly fast real-time rendering speed achieved by 3DGS, our method not only maintains comparable rendering quality across all test scenarios but also achieves further improvements in rendering speed and training speed. This enhancement is attributed to the substantial reduction in the number of Gaussians facilitated by our adaptive growth and pruning strategies. Most notably, coupled with the efficient scene representation and Gaussian management of Gaussian-Forest, our method achieves a remarkable $7\sim 17$ times reduction in model size compared to 3DGS, depending on the dataset and settings. Beyond the advantages of faster rendering speed and a significantly reduced model size, our approach has remarkably exceeded the rendering quality achieved by 3DGS on Deep Blending. This thoroughly validates the effectiveness of our approach.
4.4.2 Qualitative Comparisons
In Fig. 4, we present a comprehensive comparison of rendering quality between our Gaussian-Forest and 3DGS Kerbl et al. (2023), as well as representative ray tracing-based methods, including Mip-NeRF360 Barron et al. (2022) and InstantNGP Müller et al. (2022). Across four scenes from distinct datasets, our findings reveal comparable or even superior quality in the synthesis of novel views. This achievement is coupled with the fastest rendering speed, as well as a remarkable compression of parameters exceeding tenfold and notable improvements in both training and rendering speeds compared with the current state-of-the-art 3DGS. These outcomes substantiate the efficacy of our proposed method, aligning perfectly with quantitative results.
4.4.3 Complexity Analysis
Assuming 3DGS necessitates $N$ Gaussians for scene modeling, its spatial complexity stands at $O(59N)$. In Gaussian-Forest, explicit attributes of each hybrid Gaussian are represented by 6 parameters in its corresponding leaf node, yielding a spatial complexity of $O(6N)$. Post-adaptive growth, the root and internal nodes account for about $1.5\%\sim 2.5\%$ and $25\%\sim 50\%$ of leaf nodes, respectively. Under the configuration of $\{\mathcal{D}_{R},\mathcal{D}_{I}\}=\{24,16\}$ and 16-bit half-float feature format, the spatial complexity of non-leaf nodes ranges from $O(2.2N)$ to $O(4.3N)$ (excluding negligible parameters for MLPs and integer pointers in internal nodes). Furthermore, our pruning strategy effectively reduces leaf nodes by $1.5\sim 3$ times and non-leaf nodes by about $1.5$ times. In the end, the overall space complexity ranges from $O(3.5N)$ to $O(7N)$, yielding a compression factor of approximately $8\sim 17$, aligning seamlessly with the quantitative results. Such a reduction in Gaussian count has also accelerated both training and rendering speed, completely offsetting the time complexity introduced by the inclusion of MLPs.
4.5 Ablation Study
We conducted ablation studies on Deep Blending (DB) scenes Hedman et al. (2018) to validate the Gaussian-Forest components: hybrid Gaussian representation, forest management, adaptive growth, and pruning strategy. Hyperparameter impact was also investigated, including growth and pruning thresholds, and feature dimensions of non-leaf nodes. The following baselines were set:
@@ 4.5 Ablation Study
+Hybrid: Employment of hybrid Gaussian defined in Eq. (6), with a trivial feature association by storing features in a hash table Müller et al. (2022) for coordinate-based lookups. We adjusted the table size $T$ to control the model’s capabilities with $T=\{18,19,20,21,22,23\}$.
+Forest: Management of hybrid Gaussians in a forest as defined in Eq. (3), with non-growing root and internal nodes. To investigate the impact of feature dimensions, we conducted three sets of experiments, labeled A, B, and C, with $\{\mathcal{D}_{R},\mathcal{D}_{I}\}$ = $\{16,8\}$, $\{24,16\}$, and $\{32,24\}$, respectively.
+Growth: Management of hybrid Gaussians in a forest as defined in Eq. (3), with adaptive growth of non-leaf nodes. To quantitatively illustrate the impacts of growth thresholds $\{\mathcal{T}_{l}\}$, we delineated four configurations, labeled A, B, C, and D, with $\{\mathcal{T}_{l}\}$ = {10, 5, 2}, {10, 2.5, 2}, {5, 5, 2}, and {5, 2.5, 2}, respectively (in $10^{-4}$ units).
@@
The results in Fig. 5 show that using hybrid Gaussians significantly reduces space requirements. However, trivial feature association negatively affects scene representation capability and rendering quality. Managing hybrid Gaussians within a static forest structure has limited superiority due to constraints imposed by relying on initialized feature nodes. Incorporating adaptive growth enables the Gaussian forest to rapidly model complex areas, enhancing scene modeling and rendering quality while keeping a compact model size. Adaptive pruning identifies and eliminates redundant Gaussians, further reducing parameters without compromising rendering accuracy. In conclusion, these findings underscore the effectiveness of our proposed Gaussian-Forest.
@@
5 Conclusion
In this paper, we presents a solid solution to the storage issues associated with 3DGS in the context of compressed scene modeling. The introduced Gaussian-Forest, with its hierarchical hybrid representation, effectively organizes 3D Gaussians into a forest structure, optimizing parameterization and addressing storage constraints. The incorporation of adaptive growth and pruning strategies ensures detailed scene representation in intricate areas while substantially reducing the overall number of Gaussians. Through extensive experiments, we demonstrate that Gaussian-Forest maintains rendering speed and quality comparable to 3DGS, while achieving an impressive compression rate exceeding 10 times.

### v2

@@ IV-A5 Features and Decoders
The two MLPs, $\mathcal{F}_{\text{rgb}}$ and $\mathcal{F}_{\text{cov}}$, are implemented using the fast fully-fused-MLPs from Tiny-CUDA-NN. Each MLP consists of 2 hidden layers and is 64 neurons wide. All features are represented in 16-bit half-float, aligning with the output of the fully-fused-MLPs.
IV-B Comparative Methods
@@
3DGS [2] stands out for its SOTA performance in rendering speed and quality, albeit with a substantial model parameter count. We primarily compare our approach with 3DGS [2], as we aim to preserve or even enhance the rendering speed and quality while reducing the parameter count. Additionally, we compared with three representative ray tracing-based radiance field methods, each employing a different type of scene representation: Plenoxels [20] with explicit scene representation; InstantNGP [24] with a hybrid representation; and Mip-NeRF360 [28] with an implicit one. These three approaches represent typical examples of different scene representation methods. Contrasting with them allows for a comprehensive demonstration of the characteristics of our method.
IV-C Datasets and Metrics
@@
Following 3DGS, our model has been evaluated across 21 diverse scenarios. Of these, 13 scenes are based on real-world captures, including all nine scenes introduced by Mip-NeRF360 [28], two scenes from the Tanks&Temples dataset [35], and two from Deep Blending [36]. Additionally, all eight synthetic scenes from the Synthetic Blender dataset [1] are incorporated. These datasets encompass large-scale unbounded outdoor environments, indoor settings, and object-centric scenes. We employ commonly used PSNR, SSIM [37], and LPIPS [38] for evaluating rendering quality. Moreover, we provide information on rendering speed, model size, and training time, along with the speed-to-size ratio for a comprehensive and straightforward comparison.
@@
IV-D Results and Analyses
IV-D1 Quantitative Comparisons
Quantitative results across four benchmarks are presented in Table I, accompanied by additional visualizations for comparisons on Mip-NeRF360 and Tanks&Temples showcased in Fig. 4. Firstly, our approach excels in adeptly balancing rendering speed and model size. Across all scenarios, GaussianForest achieves the highest speed-to-size ratio, surpassing all comparative methods by a large margin while ensuring high-fidelity rendering quality.
In addition, compared to the unprecedentedly fast real-time rendering speed achieved by 3DGS, our method not only maintains comparable rendering quality across all test scenarios but also achieves further improvements in rendering speed and training speed. This enhancement is attributed to the substantial reduction in the number of Gaussians facilitated by our adaptive growth and pruning strategies. Most notably, coupled with the efficient scene representation and Gaussian management of GaussianForest, our method achieves a remarkable $7\sim 17$ times reduction in model size compared to 3DGS, depending on the dataset and settings. Beyond the advantages of faster rendering speed and a significantly reduced model size, our approach has remarkably exceeded the rendering quality achieved by 3DGS on Deep Blending. This thoroughly validates the effectiveness of our approach.
IV-D2 Qualitative Comparisons
In Fig. 5, we present a comprehensive comparison of rendering quality between GaussianForest and 3DGS [2], as well as representative ray tracing-based methods, including Mip-NeRF360 [28] and InstantNGP [24]. Across distinct datasets, our findings reveal comparable or even superior quality in the synthesis of novel views. This achievement is coupled with the fastest rendering speed, as well as a remarkable compression of parameters exceeding tenfold and notable improvements in both training and rendering speeds compared with the current state-of-the-art 3DGS. These outcomes substantiate the efficacy of our proposed method, aligning closely with quantitative results.
IV-D3 Complexity Analysis
Assuming 3DGS necessitates $N$ Gaussians for scene modeling, its spatial complexity stands at $O(59N)$. In GaussianForest, explicit attributes of each hybrid Gaussian are represented by 6 parameters in its corresponding leaf node, yielding a spatial complexity of $O(6N)$. Post-adaptive growth, the root and internal nodes account for about $1.5\%\sim 2.5\%$ and $25\%\sim 50\%$ of leaf nodes, respectively. Under the configuration of $\{\mathcal{D}_{R},\mathcal{D}_{I}\}=\{24,16\}$ and 16-bit half-float feature format, the spatial complexity of non-leaf nodes ranges from $O(2.2N)$ to $O(4.3N)$ (excluding negligible parameters for MLPs and integer pointers in internal nodes). Furthermore, our pruning strategy effectively reduces leaf nodes by $1.5\sim 3$ times and non-leaf nodes by about $1.5$ times. In the end, the overall space complexity ranges from $O(3.5N)$ to $O(7N)$, yielding a compression factor of approximately $8\sim 17$, aligning seamlessly with the quantitative results. Such a reduction in Gaussian count has also accelerated both training and rendering speed, completely offsetting the time complexity introduced by the inclusion of MLPs.
V Ablation Study
Ablation studies were conducted on Deep Blending scenes [36] to validate the core components of GaussianForest, including hybrid representation, forest management, and adaptive growth and pruning strategies. We also investigated the impact of various hyperparameters, such as growth and pruning thresholds and the feature dimensions of non-leaf nodes.
V-A Baseline Setup
V-A1 +Hybrid
This baseline employs the hybrid Gaussian representation defined in Eq. (6). Diverging from the management of explicit and implicit attributes via a forest structure, this baseline adopts a straightforward feature association approach, connecting each hybrid Gaussian to its corresponding feature based on its position. In particular, spatial features are retained within a multi-resolution hash table $\mathbf{H}$ as firstly introduced in [24], with the configuration adhering to its default settings. The corresponding feature $\mathbf{f}$ of the hybrid Gaussian $\mathbf{\Theta}_{\text{GF}}$ is retrieved by indexing $\mathbf{H}$ with its position $\bm{\mu}$, denoted by $\mathbf{f}=\mathbf{H}_{[\bm{\mu}]}$. We adjusted the table size $T$ to control the model’s capabilities with $T=\{18,19,20,21,22,23\}$.
V-A2 +Forest
This baseline integrates the hybrid Gaussian representation defined in Eq. (6), akin to +Hybrid. However, instead of the hashing feature association approach, +Forest organizes these hybrid representations within a forest structure as defined in Eq. (3). In addition, +Forest follows the initialization configuration of GaussianForest described in Sec. III-C1, but does not undergo the adaptive growth and pruning procedure. Specifically, a forest composed of trees with $L=3$ layers is adopted, with the number of nodes in both the root and internal layers initialized to $K=10$k. No nodes are added or removed during the training process. To investigate the impact of feature dimensions $\mathcal{D}_{R}$ and $\mathcal{D}_{I}$ on the model’s capabilities, we conducted three sets of experiments:
  • +ForestA: $\{\mathcal{D}_{R},\mathcal{D}_{I}\}=$16, 8}
  • +ForestB: $\{\mathcal{D}_{R},\mathcal{D}_{I}\}=$24, 16}
  • +ForestC: $\{\mathcal{D}_{R},\mathcal{D}_{I}\}=$32, 24}
V-A3 +Growth
This baseline builds upon +ForestA and additionally incorporates adaptive growth of non-leaf nodes defined in Sect. III-C2. The only distinction between +Growth and our comprehensive GaussianForest model lies in the exclusion of the pruning procedure. Moreover, we delineated four configurations for the growth thresholds $\{\mathcal{T}_{l}\}$ to quantitatively illustrate their impacts. Specifically, $\mathcal{T}_{2}$ is consistently set to 2, aligning with the Gaussian densification strategy employed in 3DGS [2], while $\mathcal{T}_{0}$ and $\mathcal{T}_{1}$ are empirically chosen from the geometric series values of 10, 5 and 2.5 (all in ×$\times 10^{-4}$ unit):
  • +GrowthA: $\{\mathcal{T}_{l}\}=$10, 5, 2}
  • +GrowthB: $\{\mathcal{T}_{l}\}=$10, 2.5, 2}
  • +GrowthC: $\{\mathcal{T}_{l}\}=$5, 5, 2}
  • +GrowthD: $\{\mathcal{T}_{l}\}=$5, 2.5, 2}
V-A4 +Pruning
@@ V-A Baseline Setup
V-B Results and Analyses
@@
Fig. 6 shows that +Hybrid, relying on position-hashing based feature association, achieves a certain level of model compression at a significant cost in rendering quality and speed compared to 3DGS [2], regardless of the hash table size setting. By adopting a hierarchical management approach and organizing hybrid Gaussians in a forest structure, +Forest notably enhances rendering quality. This highlights the potential of hierarchical-hybrid 3D Gaussian representation for compressed scene modeling.
Nevertheless, the representational capacity of the Gaussian Forest without adaptive growth of feature nodes is severely constrained by the initialization process and feature dimensions. Specifically, leaf nodes initially assigned to the same parent node remain bound together, regardless of how many times they have been cloned. Eventually, the number of leaf nodes may reach millions, while the number of feature nodes remains fixed at the initial $K=10$k. It is evident that the expressive demands cannot be met when hundreds of leaf nodes share a single feature vector in their common parent.
Furthermore, it can be observed that +ForestB, compared to +ForestA, significantly improves rendering quality by increasing the feature dimensions. Although +ForestC further enhances quality, there is a trade-off with decreased rendering speed and memory efficiency. Therefore, we choose +ForestB as our primary model configuration (Small) and the basis for subsequent ablation studies, reserving +ForestC for the Large setting when pursuing higher quality.
It can be observed that by allowing feature nodes to adaptively grow directed by cumulative gradients, +Growth achieves a substantial improvement in rendering quality with a slight increase in parameter count compared to +Forest. Notably, its rendering quality even surpasses that of 3DGS [2], a model nearly ten times larger than ours. This outcome underscores the rationality of our motivation and the effectiveness of our design, i.e., the hierarchical growth of branches in areas characterized by under-reconstruction or high uncertainty. This strategy enhances the model’s ability to depict intricate areas with finer detail through increased feature dimensions.
Moreover, as the growth threshold gradually decreases from +GrowthA to +GrowthD, there is an augmentation in the number of feature nodes, coupled with heightened rendering quality, an expanded model size, and a reduction in rendering speed. Given the marginal disparity in rendering quality and the trade-off of a smaller parameter count and faster rendering speed between +GrowthC and +GrowthD, we have chosen +GrowthC as our default parameter setting and the foundation for subsequent ablation studies.
From Fig. 6, it is evident that introducing the pruning strategy markedly boosts the rendering speed. Given the fact that the total count of Gaussians critically determines the rendering speed by influencing sorting in rasterization and $\alpha$-blending in shading, this advancement is attributed to the effective reduction of the total number of Gaussians compared with +Growth and [2] (approximately half or two-thirds). Moreover, mild pruning not only retains but can slightly improve rendering quality. This is because removed Gaussians often contribute insignificantly to the rendering process and scene representation due to their low transparency and scale. Additionally, their removal induces a subtle re-adjustment in the remaining nearby Gaussians, enabling them to compensate for the minor impacts of this elimination through ongoing training. Ultimately, in light of the balance between rendering quality and time-space efficiency, we have chosen +PruningE as our definitive GaussianForest model configuration.
VI Conclusion
In this paper, we presents a solid solution to the storage issues associated with 3DGS in the context of compressed scene modeling. The introduced GaussianForest, with its hierarchical hybrid representation, effectively organizes 3D Gaussians into a forest structure, optimizing parameterization and addressing storage constraints. The incorporation of adaptive growth and pruning strategies ensures detailed scene representation in intricate areas while substantially reducing the overall number of Gaussians. Through extensive experiments, we demonstrate that GaussianForest maintains rendering speed and quality comparable to 3DGS, while achieving an impressive compression rate exceeding 10 times.